# Weighted Chernoff information and optimal loss exponent in context-sensitive hypothesis testing

### 09 Mar 2026 v1

@@ 4 Examples and applications
The identities of Section 3 reduce the computation of $D^{\mathrm{w}}_{B,\alpha}$ and $D^{\mathrm{w}}_{C}$ to evaluating the single-letter weighted affinity $\rho^{\mathrm{w}}_{\alpha}(p,q)=\int\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x)$ (and then optimising over $\alpha\in[0,1]$ when computing $D^{\mathrm{w}}_{C}$). We work this out for several standard parametric families (Gaussian, Poisson, and exponential), highlighting how the context weight $\varphi$ modifies the classical formulas. When $\varphi\equiv 1$, the expressions reduce to the usual (unweighted) Bhattacharyya/Chernoff quantities available in the literature (e.g. [7, 9]). More involved computations outside the exponential-family setting (such as the Cauchy location–scale family) are deferred to the appendices.
@@ 4.1 Gaussian models
$=\frac{|\Sigma_{\alpha}|^{1/2}}{|\Sigma_{1}|^{\alpha/2}|\Sigma_{2}|^{(1-\alpha)/2}}\exp\!\left\{-\frac{1}{2}\Big(\alpha\mu_{1}^{\mathrm{T}}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\mathrm{T}}\Sigma_{2}^{-1}\mu_{2}-\widetilde{\mu}_{\alpha}^{\mathrm{T}}\Sigma_{\alpha}^{-1}\widetilde{\mu}_{\alpha}\Big)\right\}.$ (4.1)
@@
$=\frac{1}{2}\left(\alpha\mu_{1}^{\mathrm{T}}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\mathrm{T}}\Sigma_{2}^{-1}\mu_{2}-\widetilde{\mu}_{\alpha}^{\mathrm{T}}\Sigma_{\alpha}^{-1}\widetilde{\mu}_{\alpha}+\ln\frac{|\Sigma_{1}|^{\alpha}|\Sigma_{2}|^{1-\alpha}}{|\Sigma_{\alpha}|}\right).$ (4.2)
In particular, setting $\gamma=0$ (i.e. $\varphi\equiv 1$) reduces (4.1) to the classical (unweighted) Gaussian Bhattacharyya distance; see, e.g., [7].
@@
$=\exp\!\left\{-\frac{\alpha(1-\alpha)}{2}\|\delta\|_{\Sigma^{-1}}^{2}+\gamma^{\mathrm{T}}\mu_{\alpha}+\frac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma\right\},$ (4.3)
@@
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\frac{\alpha(1-\alpha)}{2}\|\delta\|_{\Sigma^{-1}}^{2}-\gamma^{\mathrm{T}}\mu_{\alpha}-\frac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma.$ (4.4)
@@
In particular, for $\gamma=0$ (i.e. $\varphi\equiv 1$) we recover $\alpha^{*}=1/2$ and $D_{C}(\mathbb{P},\mathbb{Q})=\|\delta\|_{\Sigma^{-1}}^{2}/8$.
@@
The expression for $\rho^{\mathrm{w}}_{\alpha}$ follows by simplifying Example 4.1 under $\Sigma_{1}=\Sigma_{2}=\Sigma$, which makes the determinant prefactor equal to $1$ and yields a Gaussian MGF term $\exp(\gamma^{\mathrm{T}}\mu_{\alpha}+\tfrac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma)$. The maximiser follows by differentiating (4.4) in $\alpha$. ∎
Choosing an exponential weight $\varphi(x)=e^{\gamma^{\mathrm{T}}x}$ corresponds to exponential tilting: for a Gaussian $X\sim\mathcal{N}(\mu,\Sigma)$ this tilting keeps the covariance and shifts the mean to $\mu+\Sigma\gamma$ (with normalisation factor $\exp(\gamma^{\mathrm{T}}\mu+\tfrac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma)$), which is why the weighted affinities remain available in closed form. In particular, the optimal Chernoff parameter is no longer forced to be $\alpha^{*}=1/2$ and, as (4.4) shows, sufficiently strong tilting can push the maximiser to the boundary $\alpha^{*}\in\{0,1\}$.
@@ 4.2 Poisson models
Let ${\cal X}=\mathbb{N}_{0}=\{0,1,2,\dots\}$ and let $\mu$ be the counting measure on $\cal X$}. Fix two hypotheses $\mathbb{P}=\mathrm{Poi}(\lambda_{1})$ and $\mathbb{Q}=\mathrm{Poi}(\lambda_{2})$ with $\lambda_{1},\lambda_{2}>0$, and write $p=p_{\lambda_{1}}$ and $q=q_{\lambda_{2}}$. We still consider the exponential weight $\varphi_{\gamma}(k)=e^{\gamma k},\penalty 10000\ \gamma\in\mathbb{R}.$ (For $\gamma=0$ we recover the unweighted case $\varphi\equiv 1$.)
For $\alpha\in[0,1]$ set
@@
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ $=\sum_{k=0}^{\infty}\varphi_{\gamma}(k)\,p(k)^{\alpha}q(k)^{1-\alpha}=\exp\!\big\{-\alpha\lambda_{1}-(1-\alpha)\lambda_{2}+e^{\gamma}\lambda_{\alpha}\big\}.$ (4.5)
@@
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\alpha\lambda_{1}+(1-\alpha)\lambda_{2}-e^{\gamma}\lambda_{\alpha}.$ (4.6)
Moreover, the normalised weighted geometric mixture (Chernoff-tilted density) from (3.5) takes the form
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}\alpha^{2}}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-e^{\gamma}\lambda_{\alpha}\left(\ln\frac{\lambda_{1}}{\lambda_{2}}\right)^{2}<0,$
hence the maximiser $\alpha^{*}$ in Definition 2.4 is unique. Differentiating (4.6) yields the critical point condition
@@ 4.3 Exponential models
Moreover, the Chernoff-tilted density $(pq)_{\alpha}$ from (3.5) is again exponential:
@@
(b) Optimal Chernoff parameter. If $\lambda_{1}\neq\lambda_{2}$, then $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})$ is strictly concave on $[0,1]$, hence the maximiser $\alpha^{*}$ in Definition 2.4 is unique. Differentiating yields the critical point condition
@@
If $\lambda_{1}=\lambda_{2}=\lambda$, then $\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\lambda/(\lambda-\gamma)$ does not depend on $\alpha$, so any $\alpha\in[0,1]$ is optimal and $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\ln(\lambda-\gamma)-\ln\lambda$. Setting $\gamma=0$ (i.e. $\varphi\equiv 1$) recovers the classical unweighted expressions.
Additional example (baseline, non-exponential family).
Appendix 7 contains a closed-form illustration for the Cauchy location–scale family. Since the Cauchy family is not an exponential family, this example complements the main text by showing that, even in the unweighted baseline case $\varphi\equiv 1$, the Bhattacharyya coefficient (in particular $\rho_{1/2}$) and the Chernoff information may involve special functions (complete elliptic integrals). For nontrivial weights $\varphi$, the symmetry $\rho_{\alpha}=\rho_{1-\alpha}$ (hence $\alpha^{*}=1/2$) typically fails and a comparable closed form is not available, so we keep the baseline Cauchy computation in the appendix.
@@ 4.4 Extension to $M$-ary hypothesis testing
We can record a direct $M$-ary analogue of Theorem 3.1 for a finite family of simple hypotheses via the optimal $M$-ary pointwise loss between pairwise minima (Lemma 4.6), and each pairwise term has the exponent given by the weighted Chernoff information. Thus the $M$-ary exponent is governed by the closest pair in terms of $D_{C}^{\mathrm{w}}$.
@@
so that all pairwise weighted Chernoff informations are well-defined.
@@
$L_{n,M}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\left(\sum_{i=1}^{M}p_{i}(x_{1}^{n})-\max_{1\leq j\leq M}p_{j}(x_{1}^{n})\right)\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (4.7)
@@
Minimisation over $\delta_{n}$ is therefore pointwise in $x_{1}^{n}$ and is achieved by selecting an index maximising $p_{j}(x_{1}^{n})$, yielding (4.7). ∎
@@
For any nonnegative numbers $a_{1},\ldots,a_{M}$,
$\max_{1\leq i<j\leq M}\min(a_{i},a_{j})\ \leq\ \sum_{k=1}^{M}a_{k}-\max_{1\leq k\leq M}a_{k}\ \leq\ \sum_{1\leq i<j\leq M}\min(a_{i},a_{j}).$ (4.8)
@@
Let $a_{(1)}\geq a_{(2)}\geq\cdots\geq a_{(M)}$ be the decreasing rearrangement of $(a_{1},\ldots,a_{M})$. Then $\sum_{k}a_{k}-\max_{k}a_{k}=\sum_{r=2}^{M}a_{(r)}\geq a_{(2)}$. Moreover, $\max\limits_{i<j}\min(a_{i},a_{j})=a_{(2)}$, proving the left inequality in (4.8).
@@
Applying (4.8) pointwise to $a_{i}=p_{i}(x_{1}^{n})$, multiplying by $\varphi(x_{1}^{n})$ and integrating yields (4.9). ∎
@@
For $1\leq i<j\leq M$, let $D_{C}^{\mathrm{w}}(\mathbb{P}_{i},\mathbb{P}_{j})$ be the weighted Chernoff information as in Definition 2.4. Set
@@
$L_{n,M}^{*}=\exp\{-nC_{M}^{\mathrm{w}}+o(n)\},\qquad n\to\infty,$ (4.10)
@@
Now use the sandwich inequality(4.9). Let $(i^{*},j^{*})$ attain the minimum $C_{M}^{\mathrm{w}}=D_{C}^{\mathrm{w}}(\mathbb{P}_{i^{*}},\mathbb{P}_{j^{*}})$. From the lower bound,
@@
Taking $\frac{1}{n}\ln(\cdot)$ and letting $n\to\infty$ yields (4.10). ∎
@@
Let $w_{1},\ldots,w_{M}>0$, $\sum_{i}w_{i}=1$, and consider the Bayesian weighted total loss $L_{n,M}^{(w)}(\delta_{n})=\sum\limits_{i=1}^{M}w_{i}L_{i,n}(\delta_{n})$ with optimum $L_{n,M}^{(w)*}:=\inf_{\delta_{n}}L_{n,M}^{(w)}(\delta_{n})$. Then the exponent remains $C_{M}^{\mathrm{w}}$. Indeed, writing $w_{\min}:=\min_{i}w_{i}$ and $w_{\max}:=\max_{i}w_{i}$, for any $\delta_{n}$,

### 09 May 2026 v3

@@ 4 Examples and Applications
The identities of Section 3 reduce the computation of $D^{\mathrm{w}}_{B,\alpha}$ and $D^{\mathrm{w}}_{C}$ to the single-letter weighted affinity
$\rho^{\mathrm{w}}_{\alpha}(p,q)=\int\varphi(x)\,p(x)^{\alpha}q(x)^{1-\alpha}\,\mathrm{d}\mu(x),$
@@
followed by optimisation over $\alpha\in[0,1]$. We work this out for Gaussian, Poisson, and exponential families, highlighting how the context weight $\varphi$ modifies the classical formulas. When $\varphi\equiv 1$, the expressions reduce to the standard unweighted Bhattacharyya and Chernoff quantities; more involved non-exponential-family computations (such as the Cauchy location–scale family) are deferred to the Appendix.
4.1 A Numerical Illustration
This subsection illustrates the behaviour of $\alpha^{\ast}$ and $D_{C}^{\mathrm{w}}$ under a non-trivial factorised weight and serves as a direct numerical verification of the Bregman identities of Section 3.3: for the model below the affinity $\rho_{\alpha}^{\mathrm{w}}$ is available in closed form, and we check that closed-form evaluation and direct numerical integration agree to machine precision.
Consider the asymmetric Gaussian hypotheses
$H_{0}:\ \mathbb{P}=\mathcal{N}(\mu_{0},\sigma_{0}^{2}),\qquad H_{1}:\ \mathbb{Q}=\mathcal{N}(\mu_{1},\sigma_{1}^{2}),\qquad(\mu_{0},\sigma_{0}^{2})=(0,1),\ (\mu_{1},\sigma_{1}^{2})=(3,2),$
with a non-indicator factorised weight
$\varphi(x)\;=\;\exp\bigl(-\beta(x-x_{0})^{2}\bigr),\qquad x_{0}=0,\ \beta\geq 0.$ (54)
At $\beta=0$, one has $\varphi\equiv 1$ and the unweighted Chernoff information is recovered. For $\beta>0$, the weight concentrates near $x_{0}=\mu_{0}$; in particular, (54) is not an indicator-type weight, so the weighted problem does not reduce to the unweighted Chernoff information on a restricted domain.
The asymmetry $\sigma_{0}^{2}\neq\sigma_{1}^{2}$ and the centring of $\varphi$ at the $H_{0}$ mean are essential for the illustration. In a fully symmetric configuration ($\sigma_{0}=\sigma_{1}$, $\mu_{1}=-\mu_{0}$, $x_{0}=0$), the problem is invariant under $\alpha\leftrightarrow 1-\alpha$, so the optimum is pinned to $\alpha^{\ast}=1/2$ for every $\beta$ and the effect of the weight on the Chernoff compromise is invisible. Asymmetric hypotheses are precisely where the weighted formalism is operationally distinct from the classical one, and it is this distinction that the numerics below is designed to expose.
Writing $A(\alpha)=\alpha/\sigma_{0}^{2}+(1-\alpha)/\sigma_{1}^{2}+2\beta$, $B(\alpha)=\alpha\mu_{0}/\sigma_{0}^{2}+(1-\alpha)\mu_{1}/\sigma_{1}^{2}+2\beta x_{0}$, and $C(\alpha)=\alpha\mu_{0}^{2}/\sigma_{0}^{2}+(1-\alpha)\mu_{1}^{2}/\sigma_{1}^{2}+2\beta x_{0}^{2}$, a direct Gaussian integration yields
$\ln\rho_{\alpha}^{\mathrm{w}}(p,q)=\tfrac{1}{2}\ln\!\tfrac{2\pi}{A(\alpha)}-\tfrac{\alpha}{2}\ln(2\pi\sigma_{0}^{2})-\tfrac{1-\alpha}{2}\ln(2\pi\sigma_{1}^{2})-\tfrac{1}{2}\Bigl(C(\alpha)-\tfrac{B(\alpha)^{2}}{A(\alpha)}\Bigr),$ (55)
and maximising (55) over $\alpha\in[0,1]$ gives $\alpha^{\ast}(\beta)$ and $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})(\beta)$. Table 1 reports their values at three selected $\beta$. Direct numerical integration of $\rho_{\alpha}^{\mathrm{w}}$ from its definition agrees with (55) to machine precision on all tabulated entries, which confirms the Bregman identities of Section 3.3 numerically for this example.
Table 1: Optimal skewing parameter and weighted Chernoff information for the Gaussian example (54), obtained by maximising (55).
$\beta$ $\alpha^{\ast}(\beta)$ $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$
0 (unweighted) $0.4153$ $0.8018$
$1/16$ $0.3355$ $0.9827$
$1/4$ $0.0963$ $1.4935$
The monotone growth of $\beta\mapsto D_{C}^{\mathrm{w}}$ and the leftward shift of $\alpha^{\ast}(\beta)$ illustrate a qualitative conclusion of Section 3: localising $\varphi$ near $\mu_{0}$ preferentially retains observations that are more probable under $H_{0}$ and thereby increases the effective discrimination rate, while simultaneously moving the optimal tilting towards the $H_{0}$ side. The classical unweighted limit is recovered at $\beta=0$.
In the language of hypothesis testing, $\alpha^{\ast}$ is the parameter that balances the exponential rates of the type-I and type-II losses at the Bayes optimum: the type-I exponent equals $\alpha^{\ast}D_{C}^{\mathrm{w}}$ and the type-II exponent equals $(1-\alpha^{\ast})D_{C}^{\mathrm{w}}$ (cf. Section 3). A leftward shift of $\alpha^{\ast}$ therefore corresponds to reallocating the available exponential budget towards faster decay of the type-II loss at the expense of the type-I loss, which is the optimal response to a weight that concentrates mass in the region where $H_{0}$ is more plausible.
Data and Code Availability
A Jupyter/Colab notebook reproducing Table 1 and Figures 13, together with the direct-integration verification of (55), is archived on Zenodo [16] and mirrored on GitHub.
Figure 1: The map $\alpha\mapsto-\ln\rho_{\alpha}^{\mathrm{w}}(p,q)$ for the Gaussian hypotheses $\mathcal{N}(0,1)$, $\mathcal{N}(3,2)$ with weight (54), for $\beta\in\{0,1/16,1/4\}$. The optimum $\alpha^{\ast}$ is marked by a bullet on each curve and shifts to the left as $\beta$ increases.
Figure 2: Optimal skewing parameter $\alpha^{\ast}(\beta)$ for the Gaussian example. The dashed line marks the unweighted value $\alpha^{\ast}(0)$.
Figure 3: Weighted Chernoff information $\beta\mapsto D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})$. The dashed line marks the classical value $D_{C}$ recovered at $\beta=0$.
@@ 4.2 Gaussian Models
$=\frac{|\Sigma_{\alpha}|^{1/2}}{|\Sigma_{1}|^{\alpha/2}|\Sigma_{2}|^{(1-\alpha)/2}}\exp\!\left\{-\frac{1}{2}\Big(\alpha\mu_{1}^{\mathrm{T}}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\mathrm{T}}\Sigma_{2}^{-1}\mu_{2}-\widetilde{\mu}_{\alpha}^{\mathrm{T}}\Sigma_{\alpha}^{-1}\widetilde{\mu}_{\alpha}\Big)\right\}.$ (56)
@@
$=\frac{1}{2}\left(\alpha\mu_{1}^{\mathrm{T}}\Sigma_{1}^{-1}\mu_{1}+(1-\alpha)\mu_{2}^{\mathrm{T}}\Sigma_{2}^{-1}\mu_{2}-\widetilde{\mu}_{\alpha}^{\mathrm{T}}\Sigma_{\alpha}^{-1}\widetilde{\mu}_{\alpha}+\ln\frac{|\Sigma_{1}|^{\alpha}|\Sigma_{2}|^{1-\alpha}}{|\Sigma_{\alpha}|}\right).$ (57)
In particular, setting $\gamma=0$ (i.e., $\varphi\equiv 1$) reduces (4.1) to the classical (unweighted) Gaussian Bhattacharyya distance; see, e.g., [12].
@@
$=\exp\!\left\{-\frac{\alpha(1-\alpha)}{2}\|\delta\|_{\Sigma^{-1}}^{2}+\gamma^{\mathrm{T}}\mu_{\alpha}+\frac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma\right\},$ (58)
@@
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\frac{\alpha(1-\alpha)}{2}\|\delta\|_{\Sigma^{-1}}^{2}-\gamma^{\mathrm{T}}\mu_{\alpha}-\frac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma.$ (59)
@@
In particular, for $\gamma=0$ (i.e., $\varphi\equiv 1$) we recover $\alpha^{*}=1/2$ and $D_{C}(\mathbb{P},\mathbb{Q})=\|\delta\|_{\Sigma^{-1}}^{2}/8$.
@@
The expression for $\rho^{\mathrm{w}}_{\alpha}$ follows by simplifying Example 4.1 under $\Sigma_{1}=\Sigma_{2}=\Sigma$, which makes the determinant prefactor equal to $1$ and yields a Gaussian MGF term $\exp(\gamma^{\mathrm{T}}\mu_{\alpha}+\tfrac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma)$. The maximiser follows by differentiating (59) in $\alpha$. ∎
Choosing an exponential weight $\varphi(x)=e^{\gamma^{\mathrm{T}}x}$ corresponds to exponential tilting: for a Gaussian $X\sim\mathcal{N}(\mu,\Sigma)$, this tilting keeps the covariance and shifts the mean to $\mu+\Sigma\gamma$ (with normalisation factor $\exp(\gamma^{\mathrm{T}}\mu+\tfrac{1}{2}\gamma^{\mathrm{T}}\Sigma\gamma)$), which is why the weighted affinities remain available in closed form. In particular, the optimal Chernoff parameter is no longer forced to be $\alpha^{*}=1/2$ and, as (59) shows, sufficiently strong tilting can push the maximiser to the boundary $\alpha^{*}\in\{0,1\}$.
@@ 4.3 Poisson Models
Let ${\cal X}=\mathbb{N}_{0}=\{0,1,2,\dots\}$ and let $\mu$ be the counting measure on $\cal X$}. Fix two hypotheses $\mathbb{P}=\mathrm{Poi}(\lambda_{1})$ and $\mathbb{Q}=\mathrm{Poi}(\lambda_{2})$ with $\lambda_{1},\lambda_{2}>0$, and write $p=p_{\lambda_{1}}$ and $q=q_{\lambda_{2}}$. Throughout this subsection we work under the standing assumption of Section 2, namely, that the observations $X_{1},\ldots,X_{n}$ are i.i.d. (distributed as $\mathbb{P}$ under $H_{0}$ and as $\mathbb{Q}$ under $H_{1}$), and that the weight $\varphi$ factorises across observations (Assumption 1.1). We still consider the exponential weight $\varphi_{\gamma}(k)=e^{\gamma k},\penalty 10000\ \gamma\in\mathbb{R}.$ (For $\gamma=0$ we recover the unweighted case $\varphi\equiv 1$.) Equivalently, setting $\varepsilon:=e^{\gamma}>0$, the weight takes the form $\varphi(k)=\varepsilon^{k}$; this reparameterisation is convenient in applications where $\varepsilon\in(0,1)$ models a per-event discount factor, while $\gamma\in\mathbb{R}$ is the natural parameter for the exponential-family calculations below. The two parameterisations are equivalent.
For $\alpha\in[0,1]$, set
@@
$\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})$ $=\sum_{k=0}^{\infty}\varphi_{\gamma}(k)\,p(k)^{\alpha}q(k)^{1-\alpha}=\exp\!\big\{-\alpha\lambda_{1}-(1-\alpha)\lambda_{2}+e^{\gamma}\lambda_{\alpha}\big\}.$ (60)
@@
$D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-\ln\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\alpha\lambda_{1}+(1-\alpha)\lambda_{2}-e^{\gamma}\lambda_{\alpha}.$ (61)
Moreover, the normalised weighted geometric mixture (Chernoff-tilted density) from (18) takes the form
@@
$\frac{\mathrm{d}^{2}}{\mathrm{d}\alpha^{2}}D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})=-e^{\gamma}\lambda_{\alpha}\left(\ln\frac{\lambda_{1}}{\lambda_{2}}\right)^{2}<0;$
hence, the maximiser $\alpha^{*}$ in Definition 2.4 is unique. Differentiating (61) yields the critical point condition
@@ 4.4 Exponential Models
Moreover, the Chernoff-tilted density $(pq)_{\alpha}$ from (18) is again exponential:
@@
(b) Optimal Chernoff parameter. If $\lambda_{1}\neq\lambda_{2}$, then $\alpha\mapsto D^{\mathrm{w}}_{B,\alpha}(\mathbb{P},\mathbb{Q})$ is strictly concave on $[0,1]$; hence, the maximiser $\alpha^{*}$ in Definition 2.4 is unique. Differentiating yields the critical point condition
@@
If $\lambda_{1}=\lambda_{2}=\lambda$, then $\rho^{\mathrm{w}}_{\alpha}(\mathbb{P},\mathbb{Q})=\lambda/(\lambda-\gamma)$ does not depend on $\alpha$, so any $\alpha\in[0,1]$ is optimal and $D_{C}^{\mathrm{w}}(\mathbb{P},\mathbb{Q})=\ln(\lambda-\gamma)-\ln\lambda$. Setting $\gamma=0$ (i.e., $\varphi\equiv 1$) recovers the classical unweighted expressions.
Additional Example (Baseline, Non-Exponential Family)
Appendix A contains a closed-form illustration for the Cauchy location–scale family. Since the Cauchy family is not an exponential family, this example complements the main text by showing that, even in the unweighted baseline case $\varphi\equiv 1$, the Bhattacharyya coefficient (in particular $\rho_{1/2}$) and the Chernoff information may involve special functions (complete elliptic integrals). For nontrivial weights $\varphi$, the symmetry $\rho_{\alpha}=\rho_{1-\alpha}$ (hence $\alpha^{*}=1/2$) typically fails and a comparable closed form is not available, so we keep the baseline Cauchy computation in the appendix.
@@ 4.5 Extension to $M$-ary Hypothesis Testing
We now record the finite-$M$ analogue of Theorem 3.1. The key observation is that the optimal $M$-ary pointwise loss is squeezed between pairwise minima (Lemma 4.6), and each pairwise term has logarithmic rate given by the corresponding weighted Chernoff information. Hence the overall $M$-ary rate is determined by the closest pair in terms of $D_{C}^{\mathrm{w}}$.
@@
so that all pairwise weighted Chernoff information values are well-defined and inequality
$\max\limits_{i\not=j}\sup\limits_{\alpha\in[0,1]}\int\varphi(x)|\ln\frac{p_{i}(x)}{p_{j}(x)}|p_{i}(x)^{\alpha}p_{j}(x)^{1-\alpha}{\mathrm{d}}\mu(x)<\infty$ (62)
holds true.
@@
$L_{n,M}^{*}=\int_{{\cal X}^{n}}\varphi(x_{1}^{n})\left(\sum_{i=1}^{M}p_{i}(x_{1}^{n})-\max_{1\leq j\leq M}p_{j}(x_{1}^{n})\right)\,\mathrm{d}\mu^{\otimes n}(x_{1}^{n}),$ (63)
@@
Minimisation over $\delta_{n}$ is therefore pointwise in $x_{1}^{n}$ and is achieved by selecting an index maximising $p_{j}(x_{1}^{n})$, yielding (63). ∎
@@
For any non-negative numbers $a_{1},\ldots,a_{M}$,
$\max_{1\leq i<j\leq M}\min(a_{i},a_{j})\ \leq\ \sum_{k=1}^{M}a_{k}-\max_{1\leq k\leq M}a_{k}\ \leq\ \sum_{1\leq i<j\leq M}\min(a_{i},a_{j}).$ (64)
@@
Let $a_{(1)}\geq a_{(2)}\geq\cdots\geq a_{(M)}$ be the decreasing rearrangement of $(a_{1},\ldots,a_{M})$. Then $\sum_{k}a_{k}-\max_{k}a_{k}=\sum_{r=2}^{M}a_{(r)}\geq a_{(2)}$. Moreover, $\max\limits_{i<j}\min(a_{i},a_{j})=a_{(2)}$, proving the left inequality in (64).
@@
Applying (64) pointwise to $a_{i}=p_{i}(x_{1}^{n})$, multiplying by $\varphi(x_{1}^{n})$ and integrating yields (65). ∎
@@
For $1\leq i<j\leq M$, let $D_{C}^{\mathrm{w}}(\mathbb{P}_{i},\mathbb{P}_{j})$ be the weighted Chernoff information as in Definition 2.4, and (62) holds true. Set
@@
$L_{n,M}^{*}=\exp\{-nC_{M}^{\mathrm{w}}+o(n)\},\qquad n\to\infty,$ (66)
@@
Now use the sandwich inequality(65). Let $(i^{*},j^{*})$ attain the minimum $C_{M}^{\mathrm{w}}=D_{C}^{\mathrm{w}}(\mathbb{P}_{i^{*}},\mathbb{P}_{j^{*}})$. From the lower bound,
@@
Taking $\frac{1}{n}\ln(\cdot)$ and letting $n\to\infty$ yields (66). ∎
@@
Let $w_{1},\ldots,w_{M}>0$, $\sum_{i}w_{i}=1$, and consider the Bayesian weighted total loss $L_{n,M}^{(w)}(\delta_{n})=\sum\limits_{i=1}^{M}w_{i}L_{i,n}(\delta_{n})$ with optimum $L_{n,M}^{(w)*}:=\inf_{\delta_{n}}L_{n,M}^{(w)}(\delta_{n})$. Then, the exponent remains $C_{M}^{\mathrm{w}}$. Indeed, writing $w_{\min}:=\min_{i}w_{i}$ and $w_{\max}:=\max_{i}w_{i}$, for any $\delta_{n}$,